OpenStudy (babyslapmafro):

Could someone please show me how 6^(3/2) simplifies to 6sqrt(6)?

ganeshie8 (ganeshie8):

another way to see this :- \(\large 6^{\frac{3}{2}}\) \(\large 6^{\frac{2+1}{2}}\) \(\large 6^{1+\frac{1}{2}}\) \(\large 6^1 6^{\frac{1}{2}}\) \(\large 6 \sqrt{6}\)

OpenStudy (loser66):

this way looks more professional than mine.